# Constraining the gravitational-wave emission of core-collapse supernovae with ground-based detectors

### v1

@@ Abstract
A gravitational-wave background (GWB) arising from the superposition of numerous unresolved gravitational-wave signals has yet to be detected. Potential contributing sources to such a background include compact binary coalescences (CBCs) and core-collapse supernovae (CCSNe). In this work, we place upper limits on the gravitational-wave energy emitted by CCSNe using cross-correlation measurements made with Advanced LIGO and Advanced Virgo detectors during their third observing run (O3). Specifically, we obtain a $95\%$ credibility upper limit of $0.01~{M_{\odot}c^{2}}$ while accounting for the contribution from CBC sources to a GWB. This result improves on previous constraint obtained from initial LIGO data by approximately two orders of magnitude. We also explore the detection prospects of third-generation ground-based detectors such as the Einstein Telescope and Cosmic Explorer for both individual CCSNe events and the GWB. Our results show that single events are likely to be detected prior to the GWB.
@@ I Introduction
Since the first discovery of gravitational waves (GWs) a decade ago Abbott et al. (2016a), numerous GW events have been detected, with 218 confirmed events reported in the version 4.0 of the Gravitational-Wave Transient Catalog (GWTC-4.0) Abac et al. (2025). However, the signals detected by ground-based interferometers such as LIGO Aasi et al. (2015), Virgo Acernese et al. (2015), and KAGRA Akutsu et al. (2019) represent only a small fraction of the GW sources in the Universe. A substantial number of GW signals remain undetected, including unresolved distant compact binary coalescences (CBCs) Zhu et al. (2011a); Marassi et al. (2011); Wu et al. (2012), core-collapse supernovae (CCSNe) Buonanno et al. (2005); Sandick et al. (2006), rotating neutron stars Ferrari et al. (1999); Regimbau and de Freitas Pacheco (2001); Zhu et al. (2011b); Lasky et al. (2013), cosmic strings Kibble (1976); Sarangi and Tye (2002); Damour and Vilenkin (2005); Abbott et al. (2018), early-universe phase transitions Lopez and Freese (2015); Dev and Mazumdar (2016); Marzola et al. (2017), etc. The collective contribution from these unresolved sources constitutes a gravitational-wave background (GWB).
@@
During the final stages of stellar evolution, stars with masses exceeding $\sim 8~{M_{\odot}}$ undergo catastrophic gravitational collapse, resulting in the formation of either a neutron star or a black hole Burrows et al. (1995); Kotake et al. (2006); Janka (2012). Throughout the collapse process, GWs are expected to be emitted. In 2021, Szczepańczyk et al. Szczepanczyk et al. (2021) compiled 82 waveforms from multidimensional CCSNe simulations, with GW energies ranging from $10^{-11}$ to $10^{-7}~{M_{\odot}c^{2}}$. In addition to these numerical simulations, several phenomenological models, such as the long-lasting bar-mode instability Ott (2010) and torus fragmentation instability Piro and Pfahl (2007), predict stronger GW emission in the range of $10^{-4}$ to $10^{-2}~{M_{\odot}c^{2}}$ Abbott et al. (2016b). Despite extensive searches, no CCSN-associated GW signals have been detected to date. The main challenge might be the intrinsically weak GW emission from CCSNe, making detection feasible only for nearby events Gossan et al. (2016); Szczepanczyk et al. (2021); Abbott et al. (2021a). Current-generation detectors are primarily sensitive to CCSNe within the Milky Way, yet the Galactic CCSNe rate is estimated to be only one or two events per century Bergh and Tammann (1991); Cappellaro et al. (1993); Tammann et al. (1994); Diehl et al. (2006); Li et al. (2011); Adams et al. (2013).
@@
In 2016, LIGO conducted a search for GWs from two CCSNe events—SN 2007gr and SN 2011dh. Although no signals were detected, limits were placed on the GW emission using sine-Gaussian waveform models. Based on the known distances to these supernovae, the estimated upper limits range from $0.1~{M_{\odot}c^{2}}$ at low frequencies to $10~{M_{\odot}c^{2}}$ at frequencies above 1 kHz Abbott et al. (2016b). During the first and second observing runs of Advanced LIGO and Advanced Virgo, additional CCSNe-targeted GW searches were conducted. Though no significant signals were found, upper limits were placed on the GW energy using ad hoc harmonic waveforms windowed with Gaussian envelopes: $E_{\text{GW}}<4.27\times 10^{-4}~{M_{\odot}c^{2}}$ at 235 Hz and $E_{\text{GW}}<1.28\times 10^{-1}~{M_{\odot}c^{2}}$ at 1304 Hz Abbott et al. (2020). More recently, the LIGO–Virgo–KAGRA Collaboration reported on GW searches targeting CCSNe, including SN 2023ixf. No GW signals were detected. Using an ellipsoidal rotating proto-neutron star model, the search sensitivity reached $E_{\text{GW}}\sim 10^{-4}~{M_{\odot}}c^{2}$ for emission at 82 Hz Abac et al. (2024).
@@
Due to the incomplete understanding of the CCSNe explosion mechanisms, current physical models remain highly uncertain. Nevertheless, GWs from CCSNe are expected to lie in the frequency range from tens Hz to several kHz, and the entire cosmological CCSNe population could contribute significantly to the GWB. Zhu et al. Zhu et al. (2010) utilized upper limits on the GWB energy density from initial LIGO data to constrain the average GW emission per CCSNe, placing bounds of $\sim 1\,M_{\odot}c^{2}$. Building on this work, we use data from the third observing run (O3) of LIGO and Virgo Abbott et al. (2021b) to place upper limits on the GWB produced by CCSNe and thereby constrain the average GW energy. Furthermore, given the non-negligible contribution of CBCs to the GWB, we extend our analysis to include CBCs alongside CCSNe. Lastly, we forecast the sensitivity of third-generation GW detectors to both individual CCSNe and their collective GWB signal.
@@
The structure of this paper is as follows. In Section II, we describe the GWB model of CCSNe and CBCs. In Section III, we use the O3 cross-correlation observations to place upper limits on $E_{\text{GW}}$. We discuss the detectability of individual CCSNe events and the GWB using future GW detectors in Section IV. Finally, Section V summarizes our findings and conclusions.
II The Gravitational wave background model
@@
$F_{\rm flux}(f,z)=\frac{1}{4\pi D_{L}(z)^{2}}\frac{\mathrm{d}E_{\text{GW}}}{\mathrm{d}f_{s}}(1+z).$ (2)
Here, $D_{L}(z)=\frac{c(1+z)}{H_{0}}\int\frac{\mathrm{d}z}{E(\Omega,z)}$ is the luminosity distance, with $E(\Omega,z)=\sqrt{\Omega_{\Lambda}+\Omega_{m}(1+z)^{3}}$. Throughout this work, we adopt the cosmological parameters $H_{0}=67.4\,\mathrm{km\,s^{-1}\,Mpc^{-1}}$, and $\Omega_{m}=1-\Omega_{\Lambda}=0.315$ Aghanim et al. (2020). The term $\mathrm{d}E_{\text{GW}}/\mathrm{d}f_{s}$ represents the GW energy spectrum emitted by a single source, with $f_{s}=f(1+z)$ being the frequency in the source frame.
Following Ref. Zhu et al. (2010), we adopt a Gaussian model to describe the average energy spectrum of CCSNe:
@@
where $A$ is the normalization amplitude, $f_{\text{peak}}$ is the spectral peak frequency, and $\Delta$ denotes the bandwidth.
@@
where $\dot{\rho}_{*}(z)$ is the cosmic star formation rate (SFR) per comoving volume in units of $M_{\odot}\,\mathrm{Mpc}^{-3}\,\mathrm{yr}^{-1}$. We adopt the parameterized form of the $\dot{\rho}_{*}(z)$ from Ref. Madau and Dickinson (2014). Here, $\mathrm{d}V/\mathrm{d}z$ is the comoving volume element, and $\Phi(m)$ is the stellar initial mass function (IMF), for which we use the modified Salpeter A form Baldry and Glazebrook (2003). We assume that each CCSN event arises from a progenitor with initial mass between 8 $M_{\odot}$ and 100 $M_{\odot}$, consistent with Ref. Smartt (2009).
@@
$\Omega_{\text{CCSNe}}(f)=\;$ $\frac{8\pi G}{3c^{2}H_{0}^{3}}f\int_{0}^{z_{*}}\frac{\dot{\rho}_{*}(z)}{(1+z)E(\Omega,z)}$
@@
Here $f_{\mathrm{ref}}=25\,\mathrm{Hz}$, and $\alpha=2/3$, which is appropriate for an inspiral-dominated CBC background spectrum; see Ref. Zhu et al. (2013) for a detailed discussion.
@@ III Upper limits on the GW energy emitted by CCSNe
Table 1: Prior ranges and posterior estimates for the GWB parameters $\log_{10}\Omega_{\text{ref}}$, $\log_{10}A$, $f_{\text{peak}}$, and $\Delta$. The first column lists the parameters. The second column shows the uniform prior ranges for Prior I, whereas Prior II are the same as Prior I except that we extend the upper ends of both $f_{\text{peak}}$ and $\Delta$ to $2000\,\mathrm{Hz}$. The third and fourth columns present median estimates and the $95\%$ credible intervals for parameters of the CCSNe and CCSNe+CBC model, respectively. For posterior estimates, values outside and inside parentheses are obtained from Prior I and Prior II, respectively.
@@
$\log_{10}\Omega_{\text{ref}}$ Uniform[13-13, 6-6] $10.6^{+2.4}_{-2.3}$ ($10.6^{+2.4}_{-2.3}$)
@@
$f_{\text{peak}}$ [Hz] Uniform[20, 1000] $520^{+457}_{-473}$ ($1044^{+919}_{-977}$) $520^{+458}_{-475}$ ($1034^{+913}_{-961}$)
$\Delta$ [Hz] Uniform[20, 1000] $502^{+474}_{-459}$ ($988^{+959}_{-928}$) $497^{+478}_{-456}$ ($977^{+973}_{-914}$)
@@
We analyze the O3 cross-correlation spectra Abbott et al. (2023) from Advanced LIGO and Virgo. The data comprise coincident strain from the H1 (LIGO Hanford), L1 (LIGO Livingston), and V1 (Virgo) detectors, and we follow the methodology of Ref. Abbott et al. (2021b). Data segments are Hann-windowed, Fourier transformed, and inverse-noise weighted to construct the cross-correlation estimator; nonstationary intervals and instrumental lines are mitigated by time-domain gating and frequency notching. Assuming the binned cross-correlation estimator is approximately Gaussian, the likelihood function for model parameters $\bm{\Theta}$ reads
@@
where $\hat{C}_{k}\equiv\hat{C}(f_{k})$ is the measured cross-correlation estimator in the $k$-th frequency bin and $\sigma(f_{k})$ is its standard deviation. We consider two GWB models: (i) CCSNe, $\Omega_{\mathrm{M}}(f)=\Omega_{\mathrm{CCSNe}}(f)$; and (ii) CCSNe + CBC, $\Omega_{\mathrm{M}}(f)=\Omega_{\mathrm{CCSNe}}(f)+\Omega_{\mathrm{CBC}}(f)$.
@@
We explore the parameter space using nested sampling implemented via dynesty Speagle (2020) within the Bilby software package Ashton et al. (2019). To validate our analysis pipeline, we reproduced the results from Ref. Abbott et al. (2021b) using the same data set and computational setup. Our independent results show excellent agreement with fig. 4 of Ref. Abbott et al. (2021b), confirming the effectiveness of our method.
@@
In Table 1, we list the prior ranges for the model parameters. For $\log_{10}\Omega_{\text{ref}}$ in the CCSNe+CBC model, we assume a uniform prior in logarithmic space over the range [13,6][-13,-6], consistent with Ref. Abbott et al. (2021b). For the peak frequency $f_{\text{peak}}$ and bandwidth $\Delta$ of the CCSNe spectrum, we assume a uniform prior between 20 and $1000\,\mathrm{Hz}$. We also assume a uniform prior for the spectrum amplitude $\log_{10}A$, for which the prior range is set such that the GW energy $E_{\text{GW}}$ is between $10^{-11}$ and $0.5~M_{\odot}c^{2}$ with 95% prior probability. These priors, as listed in Table 1, are denoted as Prior I. In order to account for peak emission above $1000\,\mathrm{Hz}$ and ultra-broad-band spectrum, for Prior II we extend the upper ends of both $f_{\text{peak}}$ and $\Delta$ to $2000\,\mathrm{Hz}$, while keeping priors for the other two parameters unchanged.
@@
Fig. 1 shows the posterior distributions of parameters for the CCSNe model (red) and the CCSNe+CBC model (blue). In Table 1, we also list the posterior estimates (including median values and 95% credible intervals) of these model parameters for both sets of priors. It can be seen that the inclusion of the CBC contribution in the GWB has a small impact on the CCSNe model posteriors. On the other hand, the CCSNe model posteriors are dependent on the prior ranges. Fig. 2 shows the 95% credibility upper limit on $\Omega_{\rm GW}(f)$ based on the posterior predictive distribution derived from the posterior distributions of parameters for the CCSNe+CBC model. It can be seen that our current limit is about a factor of three above the predicted CBC GWB signal amplitude as informed by the GW detection catalog.
@@
Figure 2: The 95% credibility upper limit on $\Omega_{\rm GW}(f)$ for the CBC+CCSNe model under two different priors. Also shown are predictions for the GWB from CBCs informed by GWTC-3 (including contributions from binary neutron stars, binary black holes and neutron star-black hole binaries), along with current and projected sensitivity curves Abbott et al. (2021b).
@@
Our upper limits are in the upper end of predictions of some phenomenological models Abbott et al. (2016b), which predict GW emission in the range of $10^{-4}$$10^{-2}~M_{\odot}c^{2}$. Comparing our constraints with the recent observational results from SN 2023ixf Abac et al. (2024), which set an upper limit of $\sim 10^{-4}~M_{\odot}c^{2}$ at 82 Hz, our results are weaker by about two orders of magnitude. However, their constraints apply at low frequencies. When extrapolated to the kHz regime, the energy bounds may increase to $10^{-2}$$10^{-1}~M_{\odot}c^{2}$, consistent with our estimates. It should be emphasized that the limit derived in this work is the average GW energy emitted per CCSN event across the cosmic population, which is complementary to single-event emission of a specific nearby supernova.
Figure 3: Assuming a generic Gaussian emission spectrum, the required GW energy $E_{\mathrm{GW}}$ as a function of peak frequency $f_{\text{peak}}$ and spectral width $\Delta$ to achieve $(S/N)_{\mathrm{GWB}}=3$ (top row) and a detection rate ($R_{\rm det}$) of 1 CCSNe event per year (bottom row). Panels correspond to different detectors A+ (left), CE (middle), and ET (right). Red dashed lines indicate constant levels of $E_{\mathrm{GW}}$.

### v2

@@ Abstract
A gravitational-wave background (GWB) arising from the superposition of numerous unresolved gravitational-wave signals has yet to be detected. Potential contributing sources to such a background include compact binary coalescences (CBCs) and core-collapse supernovae (CCSNe). In this work, we place upper limits on the gravitational-wave energy emitted by CCSNe using cross-correlation measurements made with Advanced LIGO and Advanced Virgo detectors during their third observing run. Specifically, we obtain a $95\%$ credibility upper limit of $0.01~{M_{\odot}c^{2}}$ while accounting for the contribution from CBC sources to a GWB. This result improves on previous constraint obtained from initial LIGO data by approximately 2 orders of magnitude. We also explore the detection prospects of third-generation ground-based detectors such as the Einstein Telescope and Cosmic Explorer for both individual CCSNe events and the GWB. Our results show that single events are likely to be detected prior to the GWB.
@@ I Introduction
Since the first discovery of gravitational waves (GWs) a decade ago Abbott and others (2016b), numerous GW events have been detected, with 218 confirmed events reported in version 4.0 of the Gravitational-Wave Transient Catalog (GWTC-4.0) Abac and others (2025a). However, the signals detected by ground-based interferometers such as LIGO Aasi and others (2015), Virgo Acernese and others (2015), and KAGRA Akutsu and others (2019) represent only a small fraction of the GW sources in the Universe. A substantial number of GW signals remain undetected, including unresolved distant compact binary coalescences (CBCs) Zhu et al. (2011b); Marassi et al. (2011); Wu et al. (2012), core-collapse supernovae (CCSNe) Buonanno et al. (2005); Sandick et al. (2006), rotating neutron stars Ferrari et al. (1999); Regimbau and de Freitas Pacheco (2001); Zhu et al. (2011a); Lasky et al. (2013), cosmic strings Kibble (1976); Sarangi and Tye (2002); Damour and Vilenkin (2005); Abbott and others (2018), early-Universe phase transitions Lopez and Freese (2015); Dev and Mazumdar (2016); Marzola et al. (2017), etc. The collective contribution from these unresolved sources constitutes a gravitational-wave background (GWB).
@@
During the final stages of stellar evolution, stars with masses exceeding $\sim 8~{M_{\odot}}$ undergo catastrophic gravitational collapse, resulting in the formation of either a neutron star or a black hole Burrows et al. (1995); Kotake et al. (2006); Janka (2012). Throughout the collapse process, GWs are expected to be emitted. In 2021, Szczepanczyk and others (2021) compiled 82 waveforms from multidimensional CCSNe simulations, with GW energies ranging from $10^{-11}$ to $10^{-6}~{M_{\odot}c^{2}}$. More recently, Vartanyan et al. (2023) and Choi et al. (2024) performed three-dimensional radiation-hydrodynamic simulations of nonrotating progenitors and found emitted GW energies spanning approximately $10^{-11}$ - $10^{-8}~{M_{\odot}c^{2}}$. In addition to these numerical simulations, several phenomenological models, such as the long-lasting bar-mode instability Ott (2010) and torus fragmentation instability Piro and Pfahl (2007), predict stronger GW emission in the range of $10^{-4}$ - $10^{-2}~{M_{\odot}c^{2}}$ Abbott and others (2016a). Despite extensive searches, no CCSN-associated GW signals have been detected to date. The main challenge might be the intrinsically weak GW emission from CCSNe, making detection feasible only for nearby events Gossan et al. (2016); Szczepanczyk and others (2021); Abbott and others (2021a). Current-generation detectors are primarily sensitive to CCSNe within the Milky Way, yet the Galactic CCSNe rate is estimated to be only one or two events per century Bergh and Tammann (1991); Cappellaro et al. (1993); Tammann et al. (1994); Diehl and others (2006); Li and others (2011); Adams et al. (2013).
@@
In 2016, LIGO conducted a search for GWs from two CCSNe events—SN 2007gr and SN 2011dh. Although no signals were detected, limits were placed on the GW emission using sine-Gaussian waveform models. Based on the known distances to these supernovae, the estimated upper limits range from $0.1~{M_{\odot}c^{2}}$ at low frequencies to $10~{M_{\odot}c^{2}}$ at frequencies above 1 kHz Abbott and others (2016a). During the first and second observing runs of Advanced LIGO and Advanced Virgo, additional CCSNe-targeted GW searches were conducted. Though no significant signals were found, upper limits were placed on the GW energy using ad hoc harmonic waveforms windowed with Gaussian envelopes: $E_{\text{GW}}<4.27\times 10^{-4}~{M_{\odot}c^{2}}$ at 235 Hz and $E_{\text{GW}}<1.28\times 10^{-1}~{M_{\odot}c^{2}}$ at 1304 Hz Abbott and others (2020). More recently, the LIGO–Virgo–KAGRA Collaboration reported on GW searches targeting CCSNe, including SN 2023ixf. No GW signals were detected. Using an ellipsoidal rotating protoneutron star model, the search sensitivity reached $E_{\text{GW}}\sim 10^{-4}~{M_{\odot}}c^{2}$ for emission at 82 Hz Abac and others (2024).
@@
Bcause of the incomplete understanding of the CCSNe explosion mechanisms, current physical models remain highly uncertain. Nevertheless, GWs from CCSNe are expected to lie in the frequency range from tens of hertz to several kilohertz, and the entire cosmological CCSNe population could contribute significantly to the GWB. Zhu et al. (2010) utilized upper limits on the GWB energy density from initial LIGO data to constrain the average GW emission per CCSNe, placing bounds of $\sim 1\,M_{\odot}c^{2}$. Building on this work, we use data from the third observing run (O3) of LIGO and Virgo Abbott and others (2021b) to place upper limits on the GWB produced by CCSNe and thereby constrain the average GW energy. Furthermore, given the non-negligible contribution of CBCs to the GWB, we extend our analysis to include CBCs alongside CCSNe. Finally, we forecast the sensitivity of third-generation GW detectors to both individual CCSNe and their collective GWB signal.
@@
The structure of this paper is as follows. In Sec. II, we describe the GWB model of CCSNe and CBCs. In Sec. III, we use the O3 cross-correlation observations to place upper limits on $E_{\text{GW}}$. We discuss the detectability of individual CCSNe events and the GWB using future GW detectors in Sec. IV. Finally, Sec. V summarizes our findings and conclusions.
II The Gravitational-wave background model
@@
$F_{\rm flux}(f,z)=\frac{1}{4\pi D_{L}(z)^{2}}\frac{\mathrm{d}E_{\text{GW}}}{\mathrm{d}f_{s}}(1+z)^{2}.$ (2)
Here, $D_{L}(z)=\frac{c(1+z)}{H_{0}}\int\frac{\mathrm{d}z}{E(\Omega,z)}$ is the luminosity distance, with $E(\Omega,z)=\sqrt{\Omega_{\Lambda}+\Omega_{m}(1+z)^{3}}$. Throughout this work, we adopt the cosmological parameters $H_{0}=67.4\,\mathrm{km\,s^{-1}\,Mpc^{-1}}$, and $\Omega_{m}=1-\Omega_{\Lambda}=0.315$ Aghanim and others (2020). The term $\mathrm{d}E_{\text{GW}}/\mathrm{d}f_{s}$ represents the GW energy spectrum emitted by a single source, with $f_{s}=f(1+z)$ being the frequency in the source frame.
Following Ref. Zhu et al. (2010), we adopt a Gaussian model to describe the average energy spectrum of CCSNe,
@@
where $A$ is the normalization amplitude, $f_{\text{peak}}$ is the spectral peak frequency, and $\Delta$ denotes the bandwidth. In this work, $f_{\text{peak}}$ and $\Delta$ are source-frame parameters and are assumed to be independent of redshift.
@@
[…]
@@
$\Omega_{\text{CCSNe}}(f)=\;$ $\frac{8\pi G}{3c^{2}H_{0}^{3}}f\int_{0}^{z_{*}}\frac{\dot{\rho}_{*}(z)}{(1+z)^{2}E(\Omega,z)}$
@@
Here $f_{\mathrm{ref}}=25\,\mathrm{Hz}$, and $\alpha=2/3$, which is appropriate for an inspiral-dominated CBC background spectrum; see Ref. Zhu et al. (2013) for a detailed discussion. Although the CBC background is expected to deviate from a simple power-law model at frequencies above a few hundred hertz (due to merger and ringdown contributions), our inference on $\Omega_{\mathrm{ref}}$ is primarily informed by the low-frequency data where the detectors are most sensitive and where the power-law approximation is valid. Therefore, we consider the power-law model to be a good approximation for the CBC background in the full analysis band as specified in Sec. III.
@@ III Upper limits on the GW energy emitted by CCSNe
Table 1: Prior ranges and posterior estimates for the GWB parameters $\log_{10}\Omega_{\text{ref}}$, $\log_{10}A$, $f_{\text{peak}}$, and $\Delta$. The first column lists the parameters. The second column shows the uniform prior ranges for Prior I, whereas Prior II is the same as Prior I except that we extend the upper ends of both $f_{\text{peak}}$ and $\Delta$ to $2000\,\mathrm{Hz}$. The third and fourth columns present median estimates and the $95\%$ credible intervals for parameters of the CCSNe and CCSNe+CBC model, respectively. For posterior estimates, values outside and inside parentheses are obtained from Prior I and Prior II, respectively.
@@
$\log_{10}\Omega_{\text{ref}}$ Uniform[13-13, 6-6] $\cdots$ $10.6^{+2.4}_{-2.3}$ ($10.6^{+2.4}_{-2.3}$)
@@
$f_{\text{peak}}$ (Hz) Uniform[20, 1000] $520^{+457}_{-473}$ ($1044^{+919}_{-977}$) $520^{+458}_{-475}$ ($1034^{+913}_{-961}$)
$\Delta$ (Hz) Uniform[20, 1000] $502^{+474}_{-459}$ ($988^{+959}_{-928}$) $497^{+478}_{-456}$ ($977^{+973}_{-914}$)
@@
[…]
@@
where $\hat{C}_{k}\equiv\hat{C}(f_{k})$ is the measured cross-correlation estimator in the $k$th frequency bin and $\sigma(f_{k})$ is its standard deviation. We consider two GWB models: (i) CCSNe, $\Omega_{\mathrm{M}}(f)=\Omega_{\mathrm{CCSNe}}(f)$ and (ii) CCSNe + CBC, $\Omega_{\mathrm{M}}(f)=\Omega_{\mathrm{CCSNe}}(f)+\Omega_{\mathrm{CBC}}(f)$. During sampling, $\Omega_{\mathrm{M}}(f_{k}|\Theta)$ is computed at each likelihood evaluation by numerically calculating Eq. (5) on the analysis frequency grid $f_{k}$.
@@
We explore the parameter space using nested sampling implemented via Dynesty Speagle (2020) within the Bilby software package Ashton and others (2019). To validate our analysis pipeline, we reproduced the results from Ref. Abbott and others (2021b) using the same dataset and computational setup. Our independent results show excellent agreement with FFig. 4 of Ref. Abbott and others (2021b), confirming the effectiveness of our method (see the Appendix and Fig. 5). We note that the GWB analysis presented in Ref. Abbott and others (2021b) accounted for calibration uncertainties Whelan et al. (2014); Sun and others (2020); Romero et al. (2021) for different detector baselines at the inference stage. The documentation of public O3 data products does not indicate whether or not calibration uncertainties are already included in the cross-correlation estimator and covariance. Since we successfully reproduce results from Ref. Abbott and others (2021b), we do not explicitly account for calibration uncertainties in this work.
@@
In Table 1, we list the prior ranges for the model parameters. For $\log_{10}\Omega_{\text{ref}}$ in the CCSNe+CBC model, we assume a uniform prior in logarithmic space over the range [13,6][-13,-6], consistent with Ref. Abbott and others (2021b). For the peak frequency $f_{\text{peak}}$ and bandwidth $\Delta$ of the CCSNe spectrum, we assume a uniform prior between 20 and $1000\,\mathrm{Hz}$. We also assume a uniform prior for the spectrum amplitude $\log_{10}A$, for which the prior range is set such that the GW energy [^1] The GW energy is calculated by integrating the source-frame energy spectrum over the interval of $f_{s}\in[20,5000]\,\mathrm{Hz}$. $E_{\text{GW}}$ is between $10^{-11}$ and $0.5~M_{\odot}c^{2}$ with 95% prior probability. These priors, as listed in Table 1, are denoted as Prior I. In order to account for peak emission above $1000\,\mathrm{Hz}$ and ultrabroadband spectrum, for Prior II we extend the upper ends of both $f_{\text{peak}}$ and $\Delta$ to $2000\,\mathrm{Hz}$, while keeping priors for the other two parameters unchanged.
@@
Figure. 1 shows the posterior distributions of parameters for the CCSNe model (red) and the CCSNe+CBC model (blue). The posterior distributions are relatively broad and, for some parameters, remain close to the corresponding priors. In the nondetection regime, the main information gain is reflected in the exclusion of the high-amplitude region of the prior parameter space and is therefore best summarized through upper limits. In Table 1, we also list the posterior estimates (including median values and 95% credible intervals) of these model parameters for both sets of priors. It can be seen that the inclusion of the CBC contribution in the GWB has a small impact on the CCSNe model posteriors. As an additional check, we perform a CBC-only inference using the same data, modeling the background as a pure power law with the inspiral index fixed to $\alpha=2/3$. We obtain $\log_{10}\Omega_{\mathrm{ref}}=-10.6^{+2.4}_{-2.2}$ ($95\%$ credible interval), which is statistically consistent with the posterior on $\Omega_{\mathrm{ref}}$ obtained in the CCSNe+CBC model. On the other hand, the CCSNe model posteriors are dependent on the prior ranges. Figure. 2 shows the 95% credibility upper limit on $\Omega_{\rm GW}(f)$ based on the posterior predictive distribution derived from the posterior distributions of parameters for the CCSNe+CBC model. Specifically, we draw samples of the model parameters from the CCSNe+CBC posterior, compute the corresponding spectrum $\Omega_{\rm GW}(f)$ for each sample over the analysis band, and then construct the posterior predictive distribution of $\Omega_{\rm GW}(f)$ at each frequency bin. The plotted $95\%$ upper-limit curve is the pointwise 95th percentile of this distribution. It can be seen that our current limit is about a factor of 3 above the predicted CBC GWB signal amplitude as informed by the GW detection catalog.
@@
Figure 2: The 95% credibility upper limit on $\Omega_{\rm GW}(f)$ for the CBC+CCSNe model under two different priors. Also shown are predictions for the GWB from CBCs informed by GWTC-3 (including contributions from binary neutron stars, binary black holes, and neutron star-black hole binaries), along with current and projected sensitivity curves Abbott and others (2021b).
@@
Our upper limits are in the upper end of predictions of some phenomenological models Abbott and others (2016a), which predict GW emission in the range of $10^{-4}$$10^{-2}~M_{\odot}c^{2}$. Comparing our constraints with the recent observational results from SN 2023ixf Abac and others (2024), which set an upper limit of $\sim 10^{-4}~M_{\odot}c^{2}$ at 82 Hz, our results are weaker by about 2 orders of magnitude. However, their constraints apply at low frequencies. When extrapolated to the kilohertz regime, the energy bounds may increase to $10^{-2}$$10^{-1}~M_{\odot}c^{2}$, consistent with our estimates. It should be emphasized that the limit derived in this work is the average GW energy emitted per CCSN event across the cosmic population, which is complementary to single-event emission of a specific nearby supernova.
Figure 3: Assuming a generic Gaussian emission spectrum, the required GW energy $E_{\mathrm{GW}}$ as a function of peak frequency $f_{\text{peak}}$ and spectral width $\Delta$ to achieve $(S/N)_{\mathrm{GWB}}=3$ (top row) and a detection rate ($R_{\rm det}$) of one CCSNe event per year (bottom row). Panels correspond to different detectors A+ (left), CE (middle), and ET (right). Red dashed lines indicate constant levels of $E_{\mathrm{GW}}$.